# Language Model Distillation: A Temporal Difference Imitation Learning Perspective

### v2

@@ Introduction
The successes of large language models (LLMs) have brought transformative advancements across many NLP tasks, albeit often at the cost of substantial computational resources due to their massive size. Model distillation is a widely adopted approach to compressing these large, high-capacity models (teachers) into smaller, efficient models (students). This practice has become standard, ranging from proprietary ones (achiam2023gpt) to open-source model families (touvron2023llama; grattafiori2024llama; guo2025deepseek; bai2023qwen; yang2024qwen2).
@@
Notably, this sequential distribution matching can be interpreted as behavior cloning (BC) (pomerleau1991efficient), within the imitation learning (IL) and inverse reinforcement learning (IRL) literature (abbeel2004apprenticeship; russell1998learning; ng2000algorithms; ziebart2008maximum; ho2016generative). While there are nuanced differences between IL and IRL, we use them interchangeably.
@@
However, BC is known to suffer from compounding errors (ross2011reduction; tu2022sample), also referred as exposure bias (ranzato2015sequence) in the context of auto-regressive models. To address these issues, the IL literature offers a rich toolbox beyond BC, for example generative adversarial training (ho2016generative), occupancy matching (syed2008apprenticeship), and more. We refer readers to osa2018algorithmic for a comprehensive review. While not exhaustive, many modern IL methods (syed2008apprenticeship; ho2016generative; ziebart2008maximum) involve modeling an inverse reward function and then optimizing a policy with respect to it using RL. These RL algorithms are often based on temporal-difference (TD) learning (sutton1998reinforcement), which estimates the long-term impact of an action (or token) given a state (or sub-sequence) to mitigate the compounding errors.
Distillation may potentially benefit from the broader IL toolbox—particularly from TD learning methods. Earlier work has already demonstrated the potential of such approaches in smaller auto-regressive models. For instance, yu2017seqgan and wu2021textgail can be viewed as early applications of ho2016generative. For modern LLMs, jia2024adversarial explored TD-style IL methods for distillation, through value moment matching (abbeel2004apprenticeship; ziebart2008maximum; swamy2021moments) in particular. However, applying TD learning to LLMs is generally challenging due to their vast state-action space $\mathcal{V}^{T}$ (snell2022offline; yu2023mathcal; havrilla2023trlx).
@@
$\textsc{proj}_{p}\left(\pi\right)(a\mid s):=\begin{cases}\pi(a\mid s)/\sum_{a\in\mathcal{A}^{\star}_{p}}\pi(a\mid s),&\text{if }a\in\mathcal{A}^{\star}_{p}\\ 0,&\text{otherwise}\end{cases}$ (8)

### v4

@@ Introduction
The successes of large language models (LLMs) have brought transformative advancements across many NLP tasks, albeit often at the cost of substantial computational resources due to their massive size. Model distillation is a widely adopted approach to compressing these large, high-capacity models (teachers) into smaller, efficient models (students). This practice has become standard, ranging from proprietary ones (achiam2023gpt) to open-source model families (touvron2023llama; grattafiori2024llama; yang2024qwen2).
@@
Notably, this sequential distribution matching can be interpreted as behavior cloning (BC) (pomerleau1991efficient), within the imitation learning (IL) and inverse reinforcement learning (IRL) literature (abbeel2004apprenticeship; ng2000algorithms; ziebart2008maximum; ho2016generative). While there are nuanced differences between IL and IRL, we use them interchangeably.
@@
However, BC is known to suffer from compounding errors (ross2011reduction; tu2022sample), also referred as exposure bias (ranzato2015sequence) in the context of auto-regressive models. To address these issues, the IL literature offers a rich toolbox beyond BC, for example generative adversarial training (ho2016generative), occupancy matching (syed2008apprenticeship), and more. We refer readers to osa2018algorithmic for a comprehensive review. While not exhaustive, many modern IL methods (ho2016generative; ziebart2008maximum) involve modeling an inverse reward function and then optimizing a policy with respect to it using RL. These RL algorithms are often based on temporal-difference (TD) learning (sutton1998reinforcement), which estimates the long-term impact of an action (or token) given a state (or sub-sequence) to mitigate the compounding errors.
Distillation may potentially benefit from the broader IL toolbox—particularly from TD learning methods. Earlier work has already demonstrated the potential of such approaches in smaller auto-regressive models. For instance, yu2017seqgan and wu2021textgail can be viewed as early applications of ho2016generative. For modern LLMs, jia2024adversarial explored TD-style IL methods for distillation, through value moment matching (abbeel2004apprenticeship; ziebart2008maximum) in particular. However, applying TD learning to LLMs is generally challenging due to their vast state-action space $\mathcal{V}^{T}$ (snell2022offline; yu2023mathcal; havrilla2023trlx).
@@
$\textsc{proj}_{p}(\pi)(a\mid s):=\begin{cases}\pi(a\mid s)/\sum_{a\in\mathcal{A}^{\star}_{p}}\pi(a\mid s),\;\text{if }a\in\mathcal{A}^{\star}_{p}\\ 0,\qquad\text{otherwise}\end{cases}$ (8)
@@ Conclusion
Acknowledgment
We acknowledge support from the NSF ACCESS program for providing computational resources under Award No. CIS250042.